OpenStudy (anonymous):

Kirsten used a log to make an umbrella stand. The log was 28 in. tall and had a radius of 7.5 in. Kirsten cut a cylindrical shape, with a radius of 5.5 in., out of the center of the log. After Kirsten cuts out the cylinder, what is the volume of the remaining portion of the log? Use 3.14 to approximate pi, and express your answer to two decimal places. in3

OpenStudy (anonymous):

just to make sure does it mean the outside part that needs to be found

OpenStudy (igreen):

First find the volume of the whole log. \(V = \pi r^2 h\) Plug in what we know: \(V = (3.14)(7.5^2)(28)\ Multiply 3.14 * 7.5 * 7.5 * 28

OpenStudy (igreen):

\(V = (3.14)(7.5^2)(28)\) *

OpenStudy (anonymous):

4945.5?

OpenStudy (igreen):

Yes, now we find the volume of the inside part that was cut out. \(V = \pi r^2 h\) \(V = (3.14)(5.5^2)(28)\) Multiply 3.14 * 5.5 * 5.5 * 28

OpenStudy (anonymous):

2659.58

OpenStudy (anonymous):

i got 2285.92 as the answer for the outer part of the log(the part that is left after the center was cut out)if that is what the question is asking

OpenStudy (igreen):

Yes, now subtract 4945.5 - 2659.58

OpenStudy (anonymous):

sry if im wrong tho

OpenStudy (igreen):

@jacobrocket173 Please do not give answers, it's against OpenStudy's Code-of-Conduct. We are here to guide the asker to the answer, so we need his/her participation. http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

?????

OpenStudy (igreen):

\(\color{blue}{\text{Originally Posted by}}\) @iGreen Yes, now subtract 4945.5 - 2659.58 \(\color{blue}{\text{End of Quote}}\) @TLHernandez

OpenStudy (anonymous):

??? what?

OpenStudy (anonymous):

nvm

OpenStudy (anonymous):

if you have a prob then tell that to the people i help too

OpenStudy (anonymous):

2285.92

OpenStudy (anonymous):

thats what i got any ways

OpenStudy (anonymous):

well in order to give help you eventually give the answer, right